OpenStudy (anonymous):

The tables represent two linear functions in a system. What is the solution to this system? (1, 0) (1, 6) (8, 26) (8, –22)

OpenStudy (solomonzelman):

which tables?

Nnesha (nnesha):

tennis :D

OpenStudy (solomonzelman):

I usually use chess table(s).

OpenStudy (solomonzelman):

anyway, you got anything for us, man?

OpenStudy (solomonzelman):

TABLE 1 TABLE 2 \(\bbox[8pt, lightcyan ,border:8px solid black]{\LARGE \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr ~~~y~~ \\ \hline \scr~-4~ & \scr 26 \\ \hline \scr~-2~ & \scr 18 \\ \hline \scr~~~0~ & \scr 10 \\ \hline \scr~~~2~ & \scr 2 \\ \hline \end{array} } ~~~~~~~ \huge{ \begin{array}{| l | c | r |} \hline \scr~~~x~~ & \scr ~~~y~~ \\ \hline \scr~-4~ & \scr 14 \\ \hline \scr~-2~ & \scr 8 \\ \hline \scr~~~0~ & \scr 2 \\ \hline \scr~~~5~ & \scr -4 \\ \hline \end{array} } }\)

OpenStudy (solomonzelman):

Can you find the slope in each table?

OpenStudy (anonymous):

table 1 = -8/2 table 2 = -6/?

OpenStudy (anonymous):

not negative sorry

OpenStudy (solomonzelman):

oh the table 2 is written incorrectly, the last point is (2,-4)

OpenStudy (solomonzelman):

ok, where are you saying it is not negative?

OpenStudy (anonymous):

i put -8/2 and -6/2 and i put negative signs when its not decreasing.

OpenStudy (solomonzelman):

oh, it surely is decreasing

OpenStudy (solomonzelman):

you can see that as x is becoming a greater positive number, the y is going deeper into negative values. And this is the case in both tables.

OpenStudy (anonymous):

oh

OpenStudy (solomonzelman):

each time x goes up by 2: in the first table y goes down by 8. in the second table y goes down by 6. Therefore: in the first table the slope is -4. in the second table the slope is -3.

OpenStudy (solomonzelman):

now, what are the y-intercepts in each of the table? (Hint: look at the tables)

OpenStudy (anonymous):

0,10 & 0,2

OpenStudy (solomonzelman):

Yes.

OpenStudy (solomonzelman):

but when you write a point, put it in parenthesis please. Like: (0,10)

OpenStudy (anonymous):

yeah

OpenStudy (solomonzelman):

So you know that: equation of a line (in a y intercept form, can be written as) `y=mx+b` where m is the slope, and b is the y intercept. you have: table 1 table 2 slope (m) -4 -3 y-intercept (b) 10 2 what will your 2 equations be?

OpenStudy (anonymous):

idk

OpenStudy (solomonzelman):

y=mx+b I will do the first equation for you: y=-4x+10

OpenStudy (solomonzelman):

can you do the second one for me?

OpenStudy (anonymous):

y= -3 + 2

OpenStudy (solomonzelman):

yes, almost

OpenStudy (solomonzelman):

y=-3x+2, like this

OpenStudy (solomonzelman):

so your two equations are y=-4x+10 y=-3x+2 solve using substitution: -3x+2=-4x+10

OpenStudy (anonymous):

oh i forgot the x

OpenStudy (solomonzelman):

yup... can you solve for x?

OpenStudy (anonymous):

yeah hold for a sec so i can solve

OpenStudy (solomonzelman):

sure

OpenStudy (anonymous):

x = 8

OpenStudy (solomonzelman):

yes.

OpenStudy (solomonzelman):

knowing that 8 equals 8 can you solve for y?

OpenStudy (solomonzelman):

that x* = 8

OpenStudy (solomonzelman):

(plug in 8 instead of x into any of the 2 equations)

OpenStudy (anonymous):

y = -22

OpenStudy (solomonzelman):

yup

OpenStudy (anonymous):

ohhh thanks man now i get it i will medal you and fan you. thank you :)

OpenStudy (solomonzelman):

No need fanning, although I certainly don't mind if you do (lol). You are one of the few questioners who has a very proper background to the subject, it is a joy to work with you. tY and yw:D

Nnesha (nnesha):

lol^^ btw that tables ^^O_O :O

OpenStudy (solomonzelman):

nice tables aren't they (:D) ?

Nnesha (nnesha):

yes :( but too much work for meh i like to use draw :P:DP